# Large Language Model Recall Uncertainty is Modulated by the Fan Effect

### v1

@@ Abstract
Jesse Roberts Tennessee Tech University Vanderbilt University jtroberts@tntech.edu                        Kyle Moore Vanderbilt University kyle.a.moore@vanderbilt.edu
@@ 1 Introduction
LLMs process information in a manner that is fundamentally different from humans. The matrix multiplications, maximum inner product search, and perceptron networks may have, at some level, been inspired by the biological neuronal system. But beyond the superficial, the systems bear no similarities. In spite of algorithmic and mechanistic dissimilarity, a growing body of work suggests that by merely training on human-language data, large language models learn to exhibit latent human-like cognitive behaviors as shown in Table 1.
Understanding the cognitive behaviors acquired from language will hopefully illuminate a model of potential utility already discussed, and is moreover essential to the successful application of LLMs in human-adjacent scenarios. Generally speaking, human-like cognitive effects may serve to smooth interactions between machine and human. Alternatively, a minority of discrepancies may severely undermine the interaction.
In this paper, we survey and expand the literature that applies cognitive science inspired evaluations to LLMs to analyze, understand, and catalog their relation to human cognition. Specifically, we evaluate a number of prominent LLMs for the presence of human-like fan effects à la Anderson and Reder (1999). We show that (1) some LLMs exhibit human-like fan effects based on the typicality of categorical items learned in pre-training; (2) some LLMs exhibit human-like fan effects based on the relative frequency of items in the model context; and (3) with uncertainty mitigated, the observed fan effect is disrupted. Our results reinforce that Mistral Jiang et al. (2023) and SOLAR Kim et al. (2023) seem to exhibit nuanced and noteworthy human-like behaviors. Additionally, the demonstrated connection between typicality and the fan effect provides in-silico evidence that fan effects are a special case of typicality as found to be the case in COBWEB models Silber and Fisher (1989), and we further observe differential fan effects consistent with human experiments.
@@ 2 Background
Fan effects are found by Silber and Fisher (1989) in probabilistic categories created by COBWEB to be a special case of another phenomenon known as the typicality effect. Typicality, first formalized and identified in humans by Rosch (1975), refers to a tendency of humans to perform recall and categorization tasks quicker when prompted with a more typical member of a category than with a less typical member of a category, with level of typicality determined by how common the features of an instance of a category are among all members of the same category and among contrasting categories. That is, both an item’s intra-category similarity and its inter-category similarity affect typicality assessments. For example, given pictures of two birds, a robin and a penguin, human subject response time will be higher when answering whether the penguin is a bird than whether the robin is a bird.
@@ 2.1 Prior Work
On the other hand, work has been done that establishes the presence of typicality effects in LLMs. Bhatia and Richie (2022) investigated a variety of cognitive effects in the BERT masked language model. They found that BERT shows positive evidence of typicality effects, including limited evidence of consistency with typicality violations common to humans. Misra et al. (2021) recreates the experiments of Rosch (1975), identifying their presence across numerous categories and models. Roberts et al. (2024b) replicates Misra et al. (2021) with the addition of artificial populations to verify that the results are robust to random variations in the model, suggesting that the behaviors are strongly supported. They also find that the population standard deviations correlate with typicality, though the correlation is reversed in causal language models. Typicality effects have also been shown to be present in computer vision models by Upadhyay et al. (2022).
@@ 3 In-Pretraining (Typicality) Fan Effect
Anderson originally observed the fan effect in the response times of humans when correctly responding to questions. However, in Silber and Fisher (1989), the authors observed human-like fan effects in a COBWEB model and found they were consistent with a special case of typicality. Based on this observation and extant work regarding the presence of typicality effects in LLMs, we hypothesize that LLMs may exhibit a fan effect induced by the relative typicality of categorical items. Specifically we formulate RQ3.1.
@@
Figure 1: Left col: Predictions are made using Mistral. Right col: Predictions are made using SOLAR. Bottom row: queried item is present (w/o uncertainty). Top row: queried item is absent (with uncertainty). Fan effects are evident in the negative Pearson correlation in the natural group above the noise floor.
@@ 3.1 Methodology
All experimental trials are conducted among a systematically perturbed population formed from each base model using PopulationLM Roberts et al. (2024b) to decrease the likelihood that obtained results are anomalous. The median value is the preferred aggregation when random sampling for the purpose of estimating a true value Doerr and Sutton (2019). Therefore, the median across each base model population is taken as the group prediction.
Based on work by Rosch (1975) regarding human typicality judgments across items in ten categories, we construct lists for each of the ten categories in Figure 1 by randomly selecting half of the items in a category. Selected items are included precisely once in a comma separated list with instructional content and two in-context examples. The in-context examples are not randomly sampled and are instead consistent across all experiments. An example prompt for the category bird and the item robin would be:
Following is a list that contains a number of birds. After the list, a bird will be judged as either present or absent in the list. If the list contains the bird, answer with present. If the list does not contain the bird, answer with absent. The list of birds is: toucan, magpie, swan, flamingo, duck, goose, blackbird, pelican, woodpecker, condor, canary, ostrich, redbird, catbird, lark, parakeet, hummingbird, bluejay, bluebird, sparrow, crow, vulture, cardinal, turkey, chicken, goldfinch, wren. According to the list, magpie is present. According to the list, kingfisher is absent. According to the list, robin is
@@
All experiments used an A100 GPU Google Colab environment. Token likelihoods were obtained using a fork of the minicons Python library Misra (2022).
The values for human typicality ratings are taken from Rosch (1975) and compared to the generated model probabilities to understand how typicality, as understood from human studies, impacts model behavior when performing recall.
@@ 3.3 Discussion
RoBERTa Liu et al. (2019), GPT-2 Radford et al. (2019), Llama-2 Touvron et al. (2023), and Llama-3 Meta (2024) were equivalently evaluated but showed no significant correlation, though Llama-3 does show a similar, slight effect.
We were surprised to find the fan effect exhibited in the TAG but not the TPG. However, in retrospect this could have been anticipated based on nuanced consideration of the experiment.
@@ 3.4 Next Steps
Finally, future work should investigate human behavior in an experiment similar to the described card experiment with the objective of understanding human fan effect behavior under mitigated uncertainty.
@@ 4 In-Context Fan Effect
We investigate the presence of fan effects as originally defined in Anderson (1974) in the context of concepts composed of categorical features. This addresses the question of whether fan effects show up in concepts defined exclusively in-context. We formulate this as RQ4.1. We augment our analysis to investigate the presence of differential fan effect as described in Radvansky and Zacks (1991), providing RQ4.2.
@@ 4.1 Methodology
We closely recreate the experimental methodology of Anderson (1974), with methods similar to those described in section 3.1 for in-pretraining fan effects. The experiment uses a generated model population of size $N=50$ with median aggregation across population to determine group prediction. As before, probabilities are obtained using the canary words “present” and “absent”.
Concepts are defined as natural language facts that pair persons, in the form of occupation labels, with places. Each fact is presented as a sentence of the form “The <occupation> is in the <place>”. Features are sampled from predefined person and place lists, each of size 20. The fan value is defined as the number of concepts that contain a given feature value. For example, if three distinct concepts indicate a person is present in the place “School”, the fan value of “School” is 3.
@@
An example prompt in which the concept <Doctor, Park> is queried would be the following. Note that <Mechanic, Mall> is included in all trials and has a guaranteed fan value of 1 for both features, while <Airport, Pilot> is absent in all trials.
Following is a list that contains a number of people and the places in which they are located. After the list, a person will be judged as either present or absent in a specified place. When asked about person A in place B, if the list says that person A is in place B, answer with present. If the list does not say that person A is in place B, answer with absent. The list of people and places is: The Nurse is in the Studio. The Police Officer is in the Bank. …The Mechanic is in the Mall. According to the list, in the Mall, the Mechanic is present. According to the list, in the Airport, the Pilot is absent. According to the list, in the Park, the Doctor is
Three experimental trials are conducted for both Mistral and SOLAR on an A100 GPU Google Colab environment with randomly generated concept lists to control for ordering effects and feature combination base rates due to semantically connected features (e.g. <Priest, Church>). These experiments are run only with Mistral and SOLAR due to those being the only models with positive results in the previous experiments.
@@ 4.2 Results
Figure 3: Results of the Anderson recreation experiments on SOLAR and Mistral Top row: queried item is absent with the model predicting true absence (with uncertainty). Bottom row: queried item is present with the model predicting true presence (w/o uncertainty). Lines of best fit shown with Pearson correlations shown in Figure 4.
@@
In Mistral, we once again see an obvious negative correlation between canary probability and fan value in the TAG predictions. This is consistent with a fan effect when evaluating absence of a concept (RQ 4.1). In the TAG, we see a stronger correlation with the fan value of the person feature than with the fan value of the place feature, supporting a positive result for RQ 4.2. This is consistent with results regarding differential fan effects in Radvansky and Zacks (1991), which found that the fan effect is mediated more by the fan of a particular item than the fan of a particular location.
@@
From the experiments regarding typicality, we expect that mitigated uncertainty in the TPG may lead to disruption of the fan effect. In confirmation, among TPG items all correlations fail to achieve a significant p value for fan value and canary probability Pearson correlation, again suggesting that mitigated uncertainty disrupts the fan effect.
@@ 4.3 Next Steps
Human cognitive experiments often include a dimension of elapsed time between training and testing time when studying memory-sensitive behaviors. Future work should consider simulating this time separation in language models. Though language models do not possess a directly analogous temporal dimension, experiments could evaluate the injection of semantic noise of varying length as a potential proxy.

### v2

@@ Abstract
Jesse Roberts Tennessee Tech University jtroberts@tntech.edu                        Kyle Moore Vanderbilt University kyle.a.moore@vanderbilt.edu
@@ 1 Introduction
LLMs process information in a manner that is fundamentally different from humans. The matrix multiplications, maximum inner product search, and perceptron networks may have, at some level, been inspired by the biological neuronal system. But beyond the superficial, the systems bear no similarities. In spite of algorithmic and mechanistic dissimilarity, a growing body of work suggests that by merely training on human-language data, large language models learn to exhibit human-like cognitive behaviors as shown in Table 1.
In this paper, we survey the work applying cognitive science inspired evaluations to LLMs to analyze, understand, and catalog their relation to human cognition. We extend the existing work by providing the first investigation of human-like fan effects à la Anderson and Reder (1999) in LLMs. This effect is specifically interesting because it has a relation to the previously studied typicality effect, and it is understood to be an expression of human categorization uncertainty that has been precisely measured through response time delay.
Our results show that (1) some LLMs exhibit human-like fan effects based on the typicality of categorical items learned in pre-training; (2) some LLMs exhibit human-like fan effects based on the relative frequency of items in the model context; and (3) with uncertainty mitigated, the observed fan effect is disrupted. Of the models tested, Mistral Jiang et al. (2023) and SOLAR Kim et al. (2023) exhibit noteworthy human-like fan effects, including nuanced differential fan effects previously observed in humans Radvansky (1999).
The results have two practical implications: LLMs learn to exhibit human-like uncertainty and that uncertainty may interfere with recall tasks. Our results additionally provide in-silico evidence that the fan effect is a special case of typicality as is true in COBWEB models Silber and Fisher (1989).
Understanding the cognitive behaviors acquired from language is essential to the successful application of LLMs in human-adjacent scenarios. Generally speaking, human-like cognitive effects may serve to smooth interactions between machine and human. Alternatively, a minority of discrepancies may serve to undermine the interactions.
@@ 2 Background
Fan effects are found by Silber and Fisher (1989) in probabilistic categories created by COBWEB to be a special case of another phenomenon known as the typicality effect. This would seem to suggest that fan effects may arise as a consequence of categorization, with a potential explanation being that items closer to the categorical center are more likely to collide with other items, leading to recall uncertainty, while items further from the center are less likely to experience aliasing.
Typicality, first formalized and identified in humans by Rosch (1975), refers to a tendency of humans to perform categorization tasks quicker when prompted with a more typical member of a category than with a less typical member of a category, with level of typicality determined by how common the features of an instance of a category are among all members of the same category and among contrasting categories. That is, both an item’s intra-category similarity and its inter-category similarity affect typicality assessments. For example, given pictures of two birds, a robin and a penguin, human subject response time will be higher when answering whether the penguin is a bird than whether the robin is a bird.
@@ 2.1 Prior Work
On the other hand, work has been done that establishes the presence of typicality effects in LLMs Misra et al. (2021); Bhatia and Richie (2022); Roberts et al. (2024b) as well as vision models Upadhyay et al. (2022). Bhatia and Richie (2022) found that BERT shows evidence of typicality effects, including consistency with typicality violations common to humans. Misra et al. (2021) recreated a subset of the experiments conducted by Rosch (1975) which were used to identify typicality effects in humans, identifying typicality effects across numerous categories and models. Roberts et al. (2024b) replicated Misra et al. (2021) with PopulationLM, establishing that the effect was not eroded when studied in a population.
Roberts et al. (2024b) found that the population standard deviations tended to positively correlate with typicality in encoder-only models, though not in decoder-only models. This suggests that the uncertainty captured by LLM variance may not be analogous to human uncertainty since LLMs are overwhelmingly based on the decoder-only architecture Roberts (2024).
@@ 3 In-Pretraining (Typicality) Fan Effect
Anderson originally observed the fan effect in the response times of humans when correctly responding to questions. However, in Silber and Fisher (1989), the authors observed human-like fan effects in a COBWEB model and found they were consistent with a special case of typicality. Based on this observation and extant work regarding the presence of typicality effects in LLMs, we hypothesize that LLMs may exhibit a fan effect induced by the relative typicality of categorical items acquired from pretraining. Specifically we formulate RQ3.1.
@@
In-pretraining Fan Effect Prompt Following is a list that contains a number of birds. After the list, a bird will be judged as either present or absent in the list. If the list contains the bird, answer with present. If the list does not contain the bird, answer with absent. The list of birds is: toucan, magpie, swan, flamingo, duck, goose, blackbird, pelican, woodpecker, condor, canary, ostrich, redbird, catbird, lark, parakeet, hummingbird, bluejay, bluebird, sparrow, crow, vulture, cardinal, turkey, chicken, goldfinch, wren. According to the list, magpie is present. According to the list, kingfisher is absent. According to the list, robin is____
LLM $P(present)$ and $P(absent)$
Figure 1: Prompt to measure presence/absence belief.
@@ 3.1 Methodology
Models: All experimental trials are conducted among a systematically perturbed population formed from each base model using PopulationLM Roberts et al. (2024b) to decrease the likelihood that obtained results are anomalous. The median value is the preferred aggregation when random sampling for the purpose of estimating a true value Doerr and Sutton (2019). Therefore, the median across each base model population is taken as the group prediction.
We choose RoBERTa Liu et al. (2019), GPT-2 Radford et al. (2019), Llama-2 Touvron et al. (2023), Llama-3 Meta (2024), Mistral Jiang et al. (2023), and SOLAR Kim et al. (2023) as the base models for the experiments. RoBERTa and GPT-2 are chosen as representatives of models previously studied and found to exhibit typicality effects Roberts et al. (2024b). However, past work has found that higher order human-like behaviors may not be exhibited in smaller models Roberts et al. (2024a). We therefore include large open source LLMs (Llama-2, Llama-3, Mistral, and SOLAR) that may be more likely to exhibit more nuanced recall effects.
Data Presentation: Based on work by Rosch (1975) regarding human typicality judgments across items in ten categories, we construct lists for each of the ten categories in Figure 3 by randomly selecting half of the items in a category. Selected items are included precisely once in a comma separated list with instructional content and two in-context examples. The in-context examples are not randomly sampled and are instead consistent across all experiments.
@@
Human Comparison: The values for human typicality ratings are taken from Rosch (1975) and compared to the generated model probabilities to understand how typicality, as understood from human studies, impacts model behavior when performing recall.
Other Hardware and Software: All experiments used an A100 GPU Google Colab environment. Token likelihoods were obtained using a fork of the minicons Python library Misra (2022).
Figure 3: Left col: Predictions are made using Mistral. Right col: Predictions are made using SOLAR. Bottom row: queried item is present (w/o uncertainty). Top row: queried item is absent (with uncertainty). Fan effects are evident in the negative Pearson correlation (shown in Figure 2) in the natural group above the noise floor.
@@ 3.2 Results
The noise floor observed in both SOLAR and Mistral is an empirical observation which warrants additional consideration. From our investigation, the fan effect in LLMs is modulated by the probability magnitude. Therefore, low probability outputs induce noise in the observation of the fan effect in the model probabilities which are shown for completeness in Figure 3 but filtered in the correlation analysis shown in Figure 2.
@@ 3.3 Discussion
RoBERTa Liu et al. (2019), GPT-2 Radford et al. (2019), Llama-2 Touvron et al. (2023), and Llama-3 Meta (2024) were equivalently evaluated but showed no significant correlation, though Llama-3 does show a similar, slight effect. We additionally conducted the correlation investigation presented using the population variance in place of the token probabilities and found no significant correlations. This reinforces the possibility put forth in Roberts et al. (2024b) that decoder-only LLM variance may not capture human-like uncertainty given fan effects are understood as an expression of human uncertainty.
Interpretation: We were surprised to find the fan effect exhibited in the TAG but not the TPG. However, in retrospect this could have been anticipated based on nuanced consideration of the experiment.
@@ 3.4 Next Steps
Future work should investigate human behavior in a scenario similar to the described card experiment to understand human fan effect behavior under mitigated uncertainty.
@@ 4 In-Context Fan Effect
We investigate the presence of fan effects as originally defined in Anderson (1974) in the context of concepts composed of categorical features. This addresses the question of whether fan effects show up in concepts defined and fan values induced exclusively in-context. We formulate this as RQ4.1. We augment our analysis to investigate the presence of differential fan effect as described in Radvansky and Zacks (1991), providing RQ4.2.
@@ 4.1 Methodology
We closely recreate the experimental methodology of Anderson (1974), with methods similar to those described in section 3.1 for in-pretraining fan effects.
Models: Based on the results regarding in-pretraining fan effects, we conduct in-context fan effect experiments with populations formed from Mistral and SOLAR using PopulationLM. The experiment uses a generated model population of size $N=50$ with median aggregation across population to determine group prediction. As before, probabilities are obtained using the canary words “present” and “absent”.
Data Presentation: Concepts are defined as natural language facts that pair persons, in the form of occupation labels, with places. Each fact is presented as a sentence of the form “The <occupation> is in the <place>”. Features are sampled from predefined person and place lists, each of size 20. The fan value is defined as the number of concepts that contain a given feature value. For example, if three distinct concepts indicate a person is present in the place “School”, the fan value of “School” is 3. Concept lists are randomly generated to control for ordering effects and feature combination base rates due to semantically connected features (e.g. <Priest, Church>).
@@
An example prompt in which the concept <Doctor, Park> is shown in Figure 4. Note that <Mechanic, Mall> is included in all trials and has a guaranteed fan value of 1 for both features, while <Airport, Pilot> is absent in all trials.
In-Context Fan Effect Prompt Following is a list that contains a number of people and the places in which they are located. After the list, a person will be judged as either present or absent in a specified place. When asked about person A in place B, if the list says that person A is in place B, answer with present. If the list does not say that person A is in place B, answer with absent. The list of people and places is: The Nurse is in the Studio. The Police Officer is in the Bank. …The Mechanic is in the Mall. According to the list, in the Mall, the Mechanic is present. According to the list, in the Airport, the Pilot is absent. According to the list, in the Park, the Doctor is____
LLM $P(present)$ and $P(absent)$
Figure 4: Prompt to measure presence/absence belief.
Human Comparison: The data pairings generated are based on the data presented to humans in Anderson and Reder (1999) which were shown to illicit the fan effect in human recall.
Other Hardware and Software: All experiments are conducted on an A100 GPU Google Colab environment. Token likelihoods were again obtained with a modified version of the minicons library Misra (2022).
@@ 4.2 Results
Figure 5: Results of the Anderson recreation experiments on SOLAR and Mistral Top row: queried item is absent with the model predicting true absence (with uncertainty). Bottom row: queried item is present with the model predicting true presence (w/o uncertainty). Lines of best fit are included. Pearson correlations shown in Figure 6.
@@
In Mistral, we once again see an obvious negative correlation between canary probability and fan value in the TAG predictions. This is consistent with a fan effect when evaluating absence of a concept (RQ 4.1). In the TAG, we see a stronger correlation with the fan value of the person feature than with the fan value of the place feature, supporting a positive result for RQ 4.2. This is consistent with results regarding differential fan effects in Radvansky and Zacks (1991), which found that the fan effect is mediated more by the fan of a particular object than the fan of a particular location.
@@
From the in-pretraining experiment, we expect that mitigated uncertainty in the TPG may lead to disruption of the fan effect. In confirmation, among TPG items all correlations fail to achieve a significant p value for fan value and canary probability Pearson correlation, again suggesting that mitigated uncertainty disrupts the fan effect.
@@ 4.3 Next Steps
Human cognitive experiments often include a dimension of elapsed time between training and testing time when studying memory-sensitive behaviors. Future work should consider simulating this time separation in language models. Though language models do not possess a directly analogous temporal dimension, experiments could evaluate the injection of semantic noise of varying length as a potential proxy. In fact such an experiment may suggest that time, to humans, is itself a form of semantic noise.
@@ 5 Conclusions
6 Practical Implications
Human-like uncertainty is shown to be present in Mistral and SOLAR in the form of a fan effect both when the fan value is induced in the pretraining of the model and in the context. However, just as found in Roberts et al. (2024b), the common measures of model uncertainty, variance and standard deviation, may not tend to correlate well with human uncertainty as quantified by the fan effect. This suggests that more work needs to be done to develop a human-consistent measure of LLM uncertainty.
Additionally, the fan effect should be considered when engaging LLMs in applications that require recall. The results here suggest that LLMs may have more trouble correctly evaluating the presence or absence of (1) items when the item is frequently present in the pretraining data and (2) coincident items when the base item is frequently present in the context of the model.
Acknowledgments
We appreciate the helpful recommendations of reviewers that led to improved presentation clarity and the insightful evaluation of the area chair.
Limitations
While this paper demonstrates that LLMs exhibit fan effects. It may be the case that the observed effects tend to be weak in comparison to the probability magnitude. So, it is unclear if LLMs fail in recall scenarios in manners consistent with fan effects.